# Learn With Imagination: Safe Set Guided State-wise Constrained Policy Optimization

### v4

@@ 4.2 S-3PO
To solve (9), we propose S-3PO inspired by recent trust region optimization methods Schulman et al. (2015). S-3PO uses KL divergence distance to restrict the policy search in (9) within a trust region around the most recent policy $\pi_{k}$. Moreover, S-3PO uses surrogate functions for the objective and constraints, which can be easily estimated from sample trajectories by $\pi_{k}$. Mathematically, S-3PO updates policy via solving the following optimization:
@@
Despite the complex forms, the objective and constraints in (10) can be interpreted in two steps. First, maximizing the objective (expected reward advantage) within the trust region (marked by the KL divergence constraint) theoretically guarantees the worst performance degradation. Second, constraining the current cost advantage $A_{D}^{\pi_{k}}$ based on the previous value $\mathcal{J}_{D}(\pi_{k})$ guarantees that the worst-case “imagionary” cost is non-positive at all steps as in (6). In turn, the original safety constraint (4) is satisifed.
We then show in Section 5 that S-3PO achieves (i) state-wise safety guarantee of satisfying (3), and (ii) bounded worst case performance degradation for policy update, by establishing new bounds on the difference in returns between two stochastic policies $\pi$ and $\pi^{\prime}$ for MMDPs.
@@ 4.3 Practical Implementation
The pseudocode of S-3PO is give as algorithm 1. Here we summarize two techniques that helps with S-3PO’s practical performance. (i) Weighted loss for cost value targets: A critical step in S-3PO involves fitting the cost increment value function, $V_{D}^{\pi}(\hat{s}_{t})$, which represents the maximum future cost increment relative to the highest state-wise cost observed so far. This function follows a non-increasing staircase pattern along the trajectory. Thus, we adopt a weighted loss function, $L_{weight}$, to penalize predictions that violate the non-increasing property: $L_{weight}=L(\hat{y}_{t}-y_{t})*(1+w*\text{1}[(\hat{y}_{t}-y_{t-1})>0])$, where $L$ denotes Mean Squared Error, $\hat{y}_{t}$ is the prediction, ${y}_{t}$ is the fitting target and ${w}$ is the penalty weight. (ii) Line Search scheduling: Constraints in (10) might become infeasible due to approximation errors. In this case, we perform a recovery update, enforcing the cost advantage $A_{D}^{\pi}$ to decrease from early training steps $k_{\mathrm{safe}}$ while focusing on reward improvements of $A^{\pi}$ towards the end of training, prioritizing safety first and reward performance later. Please check Section F.3 for furthur details.
@@ 5 Theoretical Results
In this section, We present three theorems, including (i) zero violation exploration (Theorem 5.1), (ii) state-wise safety guarantee of S-3PO (Theorem 5.2), and (iii) worst case performance guarantee of S-3PO (Theorem 5.3). The proofs of the three theorems are summarized in Appendix C, Appendix D and Appendix E, respectively.
Theorem 5.1 (Zero Training Time Violation).
If the system satisfies Assumption 1, and the safety index design follows the rule described in (11), the implicit safe set algorithm ensures the system is forward invariant in $\mathcal{\bar{S}}\subset\mathcal{S}_{S}$.
Theorem 5.2 (Safety Guarantee of S-3PO).
If $\pi_{k},\pi_{k+1}$ are related by applying S-3PO and $s_{t}\in\mathcal{\bar{S}}$, then $s_{t+1}=f(s_{t},\pi_{k+1}(s_{t}))\in\mathcal{\bar{S}}$ in expectation.
Theorem 5.3 (Worst-case Performance Degradation in S-3PO.).
After $k_{safe}$ updates, suppose $\pi_{k},\pi_{k+1}$ are related by S-3PO update rules, with $\epsilon^{\pi_{k+1}}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim\pi_{k+1}}[A^{\pi_{k}}(\hat{s},a)]|$, then performance return for $\pi_{k+1}$ satisfies
@@
Remark 5.4.
Theorem 5.1 shows zero safety violation exploration by proving the system state will never leave $\mathcal{\bar{S}}$ under the safeguard of ISSA. Theorem 5.2 shows that by satisfying the constraint of (10), the new policy is guaranteed to generate safe action, i.e. $a\in\mathcal{A}_{S}^{D}$, in expectation. Intuitively, Theorem 5.3 shows that with enough training epochs, the reward performance of S-3PO will not deteriorate too much after each update.
@@ 6 Experiments
In our experiments, we aim to answer: Q1: Does S-3PO achieve zero-violation during the training? Q2: How does S-3PO without safeguard compare with other state-of-the-art safe RL methods? Q3: Does S-3PO learn to act without safeguard? Q4: How does weighted loss trick impact the performance of S-3PO? Q5: Is “imaginary” cost necessary to achieve zero violation? Q6: How does S-3PO scale to high dimensional robots?
@@ 6.1 Experiments Setup
To answer these questions, we conducted our experiments on the safe reinforcement learning benchmark GUARD Zhao et al. (2023a) which is based on Mujoco and Gym interface.
@@
Comparison Group The methods in the comparison group include: (i) unconstrained RL algorithm TRPO (Schulman et al., 2015) and TRPO-ISSA. (ii) end-to-end constrained safe RL algorithms CPO (Achiam et al., 2017), TRPO-Lagrangian (Bohez et al., 2019), PCPO (Yang et al., 2020), SCPO (Zhao et al., 2023b). (iii) We select TRPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf. The full list of parameters of all methods compared can be found in Section F.2.
Metrics For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost and (iii) cost rate. More details are provided in Section F.4. We set the limit of cost to 0 for all the safe RL algorithms since we aim to avoid any violation of the constraints.
@@ 6.2 Evaluating S-3PO and Comparison Analysis
Zero Violation During Training The training performance of four representative test suites are summarized in LABEL:fig:zero_violation_performance, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Theorem 5.1. For more experiments, please check Appendix F. This superior performance is attributed to the safeguard mechanism within the S-3PO framework, which effectively corrects unsafe actions at every step, particularly during training. Furthermore, as demonstrated in LABEL:fig:comparison_results-main, the reward performance remains comparable to state-of-the-art baselines. This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, addressing Q1.
State-wise Safety Without Safety Monitor At the end of each epoch, the S-3PO policy is tested over 10,000 steps without the safeguard. This allows us to determine whether S-3PO effectively learns a state-wise safe policy through the guidance of the safe set-guided cost. As shown in in LABEL:fig:comparison_results-main, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance. These findings are consistent with Theorem 5.2 and Theorem 5.3, highlighting that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
Learn to Act without Safeguard As highlighted in 1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation. To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in LABEL:fig:comparison_results_ISSA-main. For comparison, TRPO-ISSA is included as a baseline, which relies continuously on the safeguard to maintain safe control. LABEL:fig:comparison_results_ISSA-main shows that S-3PO dramatically reduces the frequency of safeguard activations, approaching zero, indicating that a state-wise safe policy has been effectively learned, thus addressing Q3.
@@ 7 Conclusion and Future Prospectus
Appendix A Implicit Safe Set Algorithm Details
Figure 6: Notations.
A.1 Circumstances and Assumptions
In our treatment, both the robot and obstacles assume the form of point-mass circles, confined by finite collision radii. The safety criterion adopts the form $\phi_{0}=\max_{i}{\phi_{0}}_{i}$, with ${\phi_{0}}_{i}=d_{min}-d_{i}$. Here, $d_{i}$ captures the separation between the robot’s center and the $i$-th obstacle, encompassing both static and non-static entities. In this context, we introduce $z$ and $w$ to signify the relative acceleration and relative angular velocity of the robot, respectively, within the obstacle’s frame, as depicted in Figure 6. Importantly, the synthesis of the safety index for 2D collision evasion remains independent of specific dynamic models, but under the following assumption:
Assumption 1 (2D Collision Avoidance)
1) The state space is bounded, and the relative acceleration and angular velocity are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $z\in[z_{min},z_{max}]$ for $z_{min}\leq 0\leq z_{max}$; 2) For all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$; 3) The discrete-time system time step $dt\to 0$; 4) At any given time, there can at most be one obstacle becoming safety critical, such that $\phi-\eta\geq 0$ (Sparse Obstacle Environment).
The bounds in the first assumption will be directly used to synthesize $\phi$. The second assumption enables us to turn the question on whether these exists a feasible control in $\mathcal{A}_{S}^{D}$ to the question on whether there exists $z$ and $w$ to decrease $\phi$. The third assumption ensures that the discrete time approximation error is small. The last assumption enables safety index design rule applicable with multiple moving obstacles.
A.2 Safety Index Synthesis
Following the rules in (Liu and Tomizuka, 2014), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. Our goal is to choose these parameters such that $\mathcal{A}_{S}^{D}(s)$ is always nonempty. By setting $\eta=0$, the parameterization rule of safety index design rule is defined as:
$\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}\leq\frac{-z_{min}}{v_{max}}$ (11)
where $v_{max}$ is the maximum relative velocity that the robot can achieve in the obstacle frame. Note that the kinematic constraints $v_{max},z_{min}$ can be obtained by sampling the environment Zhao et al. (2021).
Appendix B Algorithms
B.1 S-3PO Algorithm
@@
        Sample nominal action $a^{r}_{t}\sim\pi_{k}(s_{t})$
        Compute and execute $a_{t}=\mathrm{ISSA}(s_{t},a^{r}_{t})$

### v5

@@ 4.3 S-3PO
To solve (9), we propose S-3PO under the framework of trust region optimization methods [Schulman et al. (2015)]. S-3PO uses KL divergence distance to restrict the policy search in (9) within a trust region around the most recent policy $\pi_{k}$. Moreover, S-3PO uses surrogate functions for the objective and constraints, which can be easily estimated from sample trajectories by $\pi_{k}$. Mathematically, S-3PO updates policy via solving the following optimization:
@@
Despite the complex forms, the objective and constraints in (10) can be interpreted in two steps. First, maximizing the objective (expected reward advantage) within the trust region (marked by the KL divergence constraint) theoretically guarantees the worst performance degradation. Second, $J_{D}(\pi)$ can not be computed at step $k+1$ since the state $s_{k+1}$ is inaccessible, thus we leverage a surrogate function to upper bound the $J_{D}(\pi)$ to guarantee the worst-case “imagionary” cost is non-positive at all steps as in (6).
@@ 4.4 Practical Implementation
The pseudocode of S-3PO is give as algorithm 1. Here we summarize two techniques that helps with S-3PO’s practical performance. (i) Weighted loss for cost value targets: A critical step in S-3PO involves fitting the cost increment value function, $V_{D}^{\pi}(\hat{s}_{t})$, which represents the maximum future cost increment relative to the highest state-wise cost observed so far. This function follows a non-increasing staircase pattern along the trajectory. Thus, we adopt a weighted loss function, $L_{weight}$, to penalize predictions that violate the non-increasing property: $L_{weight}=L(\hat{y}_{t}-y_{t})*(1+w*\text{1}[(\hat{y}_{t}-y_{t-1})>0])$, where $L$ denotes Mean Squared Error, $\hat{y}_{t}$ is the prediction, ${y}_{t}$ is the fitting target and ${w}$ is the penalty weight. (ii) Line Search scheduling: Constraints in (10) might become infeasible due to approximation errors. In this case, we perform a recovery update, enforcing the cost advantage $A_{D}^{\pi}$ to decrease from early training steps $k_{\mathrm{safe}}$ while focusing on reward improvements of $A^{\pi}$ towards the end of training, prioritizing safety first and reward performance later. Check Section C.3 for furthur details.
@@ 5 Theoretical Results
In this section, we first present the lemma to show the equivalence between constraining the imaginary cost and constraining the safety violation. Then we present the main conclusion for S3PO.
Lemma 5.1 (Safety Equivalence under Imaginary Cost).
For a given policy $\pi$ and any initially safe state $s_{0}$ ($\phi(s_{0})\leq 0$), the following two conditions are equivalent: 1) the corresponding trajectory in the training environment has zero imaginary cost $\max_{t}\Delta\phi_{t}\leq 0$; 2) the corresponding trajectory in the evaluation environment has zero safety violation $\max_{t}{\phi_{t}}\leq 0$. And if either condition holds, the two trajectories are the same.
Theorem 5.2 (Safety and Optimality of S-3PO).
S-3PO will converge in the training environment to a policy $\pi$ with no imaginary cost in expectation, and bounded worst case reward performance. In particular, if $\pi_{k}$ and $\pi_{k+1}$ are related by applying S-3PO (10), then with $\epsilon^{\pi_{k+1}}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim\pi_{k+1}}[A^{\pi_{k}}(\hat{s},a)]|$, the performance of $\pi_{k+1}$ in the training environment satisfies:
@@
Since the training environment is deterministic and all the assumptions in the theoretical results for SCPO are satisfied, the proof for the theorem directly follows from Proposition 2 from SCPO [Zhao et al. (2023b)].
By lemma 5.1, no imaginary cost in the training environment implies no safety violation in the evaluation environment. The theorem then implies that the converged policy could achieve zero safety violation in expectation in the evaluation environment. Nevertheless, formally establishing their equivalence in the probability space (e.g., in expectation) will be left for future work.
@@ 6 Experiments
In our experiments, we aim to answer: Q1: Does S-3PO achieve zero-violation during the training? Q2: How does S-3PO without safeguard compare with other advanced safe RL methods? Q3: Does S-3PO learn to act without safeguard? Q4: How does weighted loss trick impact the performance of S-3PO? Q5: Is “imaginary” cost necessary to make the RL policy learn to achieve zero violation by itself? Q6: How does S-3PO scale to high dimensional robots?
@@ 6.1 Experiments Setup
To answer these questions, we conducted experiments on the safe reinforcement learning benchmark GUARD [Zhao et al. (2023a)] which is based on Mujoco and Gym interface.
@@
Comparison Group The methods in the comparison group include: (i) unconstrained RL algorithm TRPO (Schulman et al., 2015) and TRPO-ISSA. (ii) end-to-end constrained safe RL algorithms CPO (Achiam et al., 2017), TRPO-Lagrangian (Bohez et al., 2019), PCPO (Yang et al., 2020), SCPO (Zhao et al., 2023b). (iii) We select TRPO as our baseline method since it already has safety-constrained derivatives that can be tested off-the-shelf. The full list of parameters of all methods compared can be found in Section C.2.
Metrics For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost, and (iii) cost rate. More details are provided in Section C.4. We set the limit of cost to 0 for all safe RL algorithms since no violation of the constraints is allowed.
@@ 6.2 Evaluating S-3PO and Comparison Analysis
Zero Violation During Training The training performance of four representative test suites are summarized in LABEL:fig:zero_violation_performance, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Zhao et al. (2021). For more experiments, please check Appendix C. This superior performance is attributed to the safeguard mechanism within the S-3PO framework, which effectively corrects unsafe actions at every step, particularly during training. Furthermore, as demonstrated in LABEL:fig:comparison_results-main, the reward performance remains comparable to advanced baselines. This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, addressing Q1.
State-wise Safety Without Safety Monitor At the end of each epoch, the S-3PO policy is tested over 10,000 steps without the safeguard. This allows us to determine whether S-3PO effectively learns a state-wise safe policy through the guidance of the safe set-guided cost. As shown in LABEL:fig:comparison_results-main, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance. These findings highlights that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
Learn to Act without Safeguard As highlighted in Section 4.1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation. To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in LABEL:fig:comparison_results_ISSA-main. For comparison, TRPO-ISSA is included as a baseline, which relies continuously on the safeguard to maintain safe control. LABEL:fig:comparison_results_ISSA-main shows that S-3PO dramatically reduces the frequency of safeguard activations, approaching zero, indicating that a state-wise safe policy has been effectively learned, thus addressing Q3.
@@ 7 Conclusion and Future Prospectus
Acknowledgement
This work is supported by the National Science Foundation under grant No. 2144489.
Appendix A Algorithms
A.1 S-3PO Algorithm
@@
        Sample nominal action $a_{t}\sim\pi_{k}(s_{t})$
        Compute and execute $a_{t}^{*}=\mathrm{ISSA}(s_{t},a^{r}_{t})$